OpenStudy (anonymous):

Find the most economical purchase: Golden Sun Vitamin E, 200 tablets for $12.99 or 400 tablets for $18.69

OpenStudy (anonymous):

price per tablet \[\frac{\text{cost}}{\text{number of tablets}}\]

OpenStudy (anonymous):

You want to find out what the unit price is, so divide the cost by the amount of pills in the bottle. which ever is less is the most economical.

OpenStudy (anonymous):

\(\frac{12.99}{200}\) vs \(\frac{18.69}{400}\) smallest number is the best deal you can probably do it in your head, since 400 is twice 200 but $18.69 is less that twice $12.99